OpenStudy (beccaboo333):

Connor went to New York for vacation. He spent 3 nights at a hotel and rented a car for 4 days. Jillian stayed at the same hotel, but spent 4 nights and rented a car for 5 days from the same company. If Connor paid $675 and Jillian paid $875, how much did one night at the hotel cost?

OpenStudy (anonymous):

You have 2 equations: Connor: 3x+4y=675 Jillian: 4x+5y=875 Can you rearrange one of the formulas so that you're just left with a variable (eg x= or y=)?

OpenStudy (beccaboo333):

???

OpenStudy (anonymous):

Do the equations make sense?

OpenStudy (beccaboo333):

no :/ im not good in math

OpenStudy (anonymous):

yh it does and i think its right....

OpenStudy (anonymous):

That's okay! So the whole cost of Connor's trip is going to be $675, and the total amount he pays is made up of the separate amounts he pays at each place. So we could look at his trip like this: 3 nights at the hotel X the amount he payed each night at the hotel + 4 days using a rented car X the amount he payed each day to rent the car = Total amount he payed. Does this make sense so far?

OpenStudy (beccaboo333):

not really..

OpenStudy (anonymous):

Okay! So let's just look at the time connor spends at the hotel. How many days does he spend there?

OpenStudy (beccaboo333):

3

OpenStudy (anonymous):

And he's going to pay a certain amount each day to stay at that hotel.

OpenStudy (anonymous):

So let's say that the amount he pays each day to stay at the hotel is X dollars.

OpenStudy (beccaboo333):

alright

OpenStudy (anonymous):

So if he spends 3 days at a rate of X dollars per day, we can say that the amount he payed at the hotel was 3X dollars. Does that make sense?

OpenStudy (beccaboo333):

yes

OpenStudy (anonymous):

And if he rented a car for 4 days at a rate of y dollars per day, we can say he spent 4y dollars on the car.

OpenStudy (anonymous):

Now, the total amount he payed on the trip is the amount he spent at the hotel plus the amount he spent on the car. If he spent 3x dollars at the hotel and 4y dollars on the car, we can say that 3x + 4y is the total amount he spent, which was $675. So the equation for connor's spendings is 3x + 4y = 675

OpenStudy (beccaboo333):

i think i know it now.. is it $125

OpenStudy (beccaboo333):

?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Do you understand it a little better now?

OpenStudy (beccaboo333):

thanks for the help :)

OpenStudy (anonymous):

No prob :)

OpenStudy (beccaboo333):

i have a few more questions

OpenStudy (anonymous):

Sure, what are they?